I dont understand the process,
2x-y=32
y-5x=13

Found 2 solutions by richard1234, S_Marshall:
Answer by richard1234(7193)

Add both equations (y's will cancel) to get

2x + (-5x) = 32 + 13 --> -3x = 45 --> x = -15

Replace x with -15 in either equation (I will use the first one) to get

2(-15) - y = 32 --> -30 - y = 32 --> y = -62

The solution is (x,y) = (-15,-62).

Answer by S_Marshall(1)

There are many ways to sole this problem, but I will turn these two equations into one:
2x-y=32
-y=-2x+32 -------make the equality equal y by subtracting 2x from both sides
y = 2x - 32---------------y can not be negative so multiply the whole side by -1


y-5x = 13-----------add 5x to make the eguality be in terms of y
y= 5x+13

Now both equalities equal y. Then you can equal it to each other. Its kinda like if a=b and a= c, then b=c.
5x + 13 = 2x - 32
3x + 13 = -32
3x = -45
x = -15

Then plug in x into anyone of the ORIGINAL equations. It doesn't matter which one because both will give you the same answer:
y-5x=13
y-5(-15)=13-------------two negatives make a positive
y + 75 = 13--------------subtract 75 from each side
y= -62


ans = (-15, -62)
The answer is written as coordinates because the two equations are lines and this is the point where they connect. To get a better visual, you can graph the two lines and see at which point they intersect which should be the answer above.